The cost of upgrading a section of Grand Loop Road in Yellowstone National Park is $1.7 million. Resurfacing and other maintenance are expected to cost $350,000 every 3 years. What is the capitalized cost of the road at an interest rate of 6% per year? Use manual equation and show steps
Answer #1

P = 1700000

i = 6% = 0.06

n = 3

Maintenance cost = 350000

Capitalized cost = -1700000-350000*(A/F,6%,3)/0.06

Use the following formula instead of A/F =

i/((1 + i)^n-1)

= -1700000-350000*0.3141/0.06 = -3532250